Calculate the number of sides of a regular polygon in which the exterior angle os one - fifth of the interior angle. Please show working.
Answer by stanbon(75887)

Calculate the number of sides of a regular polygon in which the exterior angle is one - fifth of the interior angle. Please show working.
----
Let the interior angle be "x"
Then the exterior angle is "x/5"
----
Equation:
x + x/5 = 180 degrees
(6/5)x = 180
(1/5)x = 30 degrees
---
Note: The sum of all the exterior angle = 360.
# of exterior angles = 360/30 = 12
---
Therefore # of sides = 12
==============
Cheers,
Stan H.